OpenStudy (anonymous):

Find the discriminant for each equation. Describe the roots. Give D, then the number of rational, irrational or complex roots. ( ex: 36, 2 rational) 25 + 4x2=-20x

OpenStudy (anonymous):

ive done this problem and i got the answer inncorrect and im not sure why just trying to compare answers with someone for my online class

OpenStudy (anonymous):

once you put the equation into polynomial form, the discrimanent is the part under the square root in the quadratic formula

OpenStudy (anonymous):

i'll work it out with you

OpenStudy (anonymous):

its (4)^2-4(25)(-20) right?

OpenStudy (anonymous):

4x^2 + 20x + 25 = 0 so a=4 b=20 c=25 then the discrimnant is b^2 - 4ac so 20^2 - 4 (4) 25 = 0 now here are the root possibilities depending on what the value of the discrimiant is: (by the way, ours is 0) if its > 0, the roots are real numbers and unequal if its = 0, the roots are real numbers and equal if its < 0, the roots are complex numbers and unequal make sense?

OpenStudy (anonymous):

oo im sorry i wrote 20 and 4 in different places :/

OpenStudy (anonymous):

could you help me on Find the discriminant for each equation. Describe the roots. Give D, then the number of rational, irrational or complex roots. ( ex: 36, 2 rational) 2x -5 = -x2

OpenStudy (anonymous):

is it (5)^2-4(2)(-1)??

OpenStudy (anonymous):

2x - 5 = -x^2 first move all the terms to one side x^2 + 2x - 5 = 0 then identify the A, B, C coeffients whcih go into the quadratic equation. now, these are the posistions of the coefficients A, B, C Ax^2 + Bx + C = 0 <- this is the standard form so we will extract A, B, and C from your equation using the placements given to us from the standard form, so A = 1 B = 2 C = -5 make sense so far?

OpenStudy (anonymous):

yuppp got it

OpenStudy (anonymous):

wait soo it would b (2)^2-4(1)(-5)

OpenStudy (anonymous):

yep, you got it!

OpenStudy (anonymous):

thank ya

OpenStudy (anonymous):

you are welcome